Carla received a 750 simple interest loan with an annual interest rate of 1.5% and a 3 year payback. How much interest will Carla pay during the 3 year term?
Answer by stanbon(75887)

Carla received a 750 simple interest loan with an annual interest rate of 1.5% and a 3 year payback. How much interest will Carla pay during the 3 year term?
-----
I = P*r*t
Interest = 750*0.015*3 = 750*0.045 = $33.75
----------
Cheers,
Stan H.
-----------